solve each system by substitution
y= -3x-20
6x+2y=5

Answer by richwmiller(17219)

y= -3x-20
6x+2y=5
6x+2(-3x-20)=5
6x+(-6x-40)=5
6x-6x-40=5
0x=45
0 =45
no solution
2y=5-6x
y=5/2 -3x
-3x-20=-3x+5/2
-20=5/2
-40=5
0=45
no solution
There is probably an error in copying the problem.